# Generating Verifiable CoT from Execution-Traces

### v1

@@ 5.2. Impact of Data Curation
For the Granite-3.3-8b model, the results are definitive. As shown in Table 2, the 25k correctness-filtered dataset substantially outperforms the others across all benchmarks. On LiveCodeBench, it achieves a score of 44.9%, an absolute gain of +26.6 over the base model, and significantly outperforms the larger 54k set (34.1%). This powerful improvement from a targeted, high-quality dataset strongly indicates that for complex reasoning, data verifiability is far more impactful than sheer volume.
When repeating the experiment with Qwen2.5-Coder-7B, all three fine-tuned models dramatically outperform the base model. For instance, the 25k dataset boosts the LiveCodeBench score from 46.3% to 67.0% (+20.7). Unlike with Granite, the performance across the three data subsets is highly competitive. The 25k dataset achieves the highest score on LiveCodeBench, while the 54k dataset has a slight edge on the CruxEval benchmarks. Given the consistently strong performance of the 25k set, we selected it as the winning configuration for all subsequent experiments.
@@ 5.3. Impact of Reasoning Direction
This finding is further confirmed with the Qwen2.5-Coder-7B model. The bi-directionally trained model once again emerges as the top performer, achieving the best results on LiveCodeBench (68.2%) and CruxEval Input (61.9% pass@1). This configuration’s final score on LiveCodeBench represents a +21.9 point gain over the base model. To contextualize this performance, Figure 7 compares our best model against a range of competitive foundation models. Our fine-tuned model establishes a leading score on LiveCodeBench-Execution and demonstrates highly competitive performance across all reasoning benchmarks.
@@
To answer this question, we systematically evaluate how the source and validation of rationales affect downstream model performance. We are not merely comparing verification strategies, but rather fundamentally different approaches to CoT generation: execution-grounded rationales (guaranteed correct by construction), LLM-generated rationales (potentially hallucinated), and no rationales at all.
@@
Table 3 presents the complete results. The ablation reveals a clear hierarchy of components, with three key insights. First, execution grounding is the dominant factor: LLM-generated CoT, despite providing step-by-step rationales, fundamentally fails because it trains models on plausible but unverified reasoning patterns. This validates our core thesis that the source of rationales—whether grounded in actual execution or generated from statistical patterns—determines whether models internalize verifiable semantics or learn superficial imitation. Second, verification quality compounds this effect: even when using execution-grounded rationales, insufficient candidate diversity degrades performance, as single solution-test pairs cannot filter edge-case failures. This confirms that statistical consensus through Dual Agreement is not merely a scalability mechanism but a necessary correctness guarantee. Third, the step-by-step reasoning itself is essential: removing CoT rationales entirely causes the largest performance drop, demonstrating that models require explicit traces of how outputs arise from inputs, not just input-output correspondences. Prompting alone provides negligible improvement, underscoring that these reasoning capabilities must be internalized through fine-tuning on verified data rather than retrieved at inference time.
@@ 5.5. Enhancing Instruction-Tuned Models
To confirm that our best-performing bi-directional data provides a specialized signal complementary to general instruction tuning, we fine-tuned existing instruct-tuned models. As shown in Figure 8, this provides a substantial boost, yielding a massive +39.9 gain on CruxEval Input for granite-3.3-8B-instruct and a +21.5 gain on CruxEval Output for the already strong Qwen2.5-Coder-7B-Instruct, demonstrating the value of our execution-grounded data.
@@ 5.6. Chain-of-Thought Consistency and Quality Analysis
To assess the quality of reasoning generated by our fine-tuned model, we compare CoTs generated by our Qwen2.5-Coder-7B-FT verified CoT model against the base Qwen2.5-Coder-7B model on CruxEval-O problems. For each problem, we generate CoT from both models and evaluate two key dimensions: consistency between reasoning and final answer, and information content of the reasoning content.

### v2

@@ 5.2. Impact of Data Curation
For the Granite-3.3-8b model, the results are definitive. As shown in Table 2, the 25k correctness-filtered dataset substantially outperforms the others across all benchmarks. On LiveCodeBench, it achieves a score of 44.9%, an absolute gain of +26.6 over the base model, and significantly outperforms the larger 54k set (34.1%). This powerful improvement from a targeted, high-quality dataset indicates that for complex reasoning, data verifiability is far more impactful than sheer volume.
When repeating the experiment with Qwen2.5-Coder-7B, all three fine-tuned models dramatically outperform the base model. For instance, the 25k dataset boosts the LiveCodeBench score from 46.3% to 67.0% (+20.7). Unlike with Granite, the performance across the three data subsets is competitive. The 25k dataset achieves the highest score on LiveCodeBench, while the 54k dataset has a slight edge on the CruxEval benchmarks. Given the consistently strong performance of the 25k set, we selected it as the winning configuration for all subsequent experiments.
@@ 5.3. Impact of Reasoning Direction
This finding is further confirmed with the Qwen2.5-Coder-7B model. The bi-directionally trained model once again emerges as the top performer, achieving the best results on LiveCodeBench (68.2%) and CruxEval Input (61.9% pass@1). For this configuration, final score on LiveCodeBench represents a +21.9 point gain over the base model. To contextualize this performance, Figure 7 compares our best model against a range of competitive foundation models. Our fine-tuned model establishes a leading score on LiveCodeBench-Execution and demonstrates highly competitive performance across all reasoning benchmarks.
@@
To answer this question, we systematically evaluate how the source and validation of rationales affect downstream model performance. We are not only comparing verification strategies, but the different approaches to CoT generation: execution-grounded rationales, LLM-generated rationales (potentially hallucinated), and no rationales at all.
@@
Table 3 presents the complete results. The ablation reveals a clear hierarchy of components, with three key insights. First, execution grounding is the dominant factor: LLM-generated CoT, despite providing step-by-step rationales, fails because it trains models on plausible but unverified reasoning patterns. This validates our core thesis that the source of rationales—whether grounded in actual execution or generated from statistical patterns—determines whether models internalize verifiable semantics or learn superficial imitation. Second, verification quality compounds this effect: even when using execution-grounded rationales, insufficient candidate diversity degrades performance, as single solution-test pairs cannot filter edge-case failures. This confirms that consensus through Dual Agreement is not merely a scalability mechanism but a necessary correctness guarantee. Third, the step-by-step reasoning itself is essential: removing CoT rationales entirely causes the largest performance drop, demonstrating that models require explicit traces of how outputs arise from inputs, not just input-output correspondences. Prompting alone provides negligible improvement, underscoring that these reasoning capabilities must be internalized through fine-tuning on verified data rather than retrieved at inference time.
@@ 5.5. Enhancing Instruction-Tuned Models
To confirm that our best-performing bi-directional data provides a specialized signal complementary to general instruction tuning, we fine-tuned existing instruct-tuned models. As shown in Figure 8, this provides a substantial boost, yielding a sharp +39.9 gain on CruxEval Input for granite-3.3-8B-instruct and a +21.5 gain on CruxEval Output for the already strong Qwen2.5-Coder-7B-Instruct, demonstrating the value of our execution-grounded data.
@@ 5.6. Chain-of-Thought Consistency and Quality Analysis
To assess the quality of reasoning generated by our fine-tuned model, we compare CoTs generated by our fine-tuned model Qwen2.5-Coder-7B-FT against the base Qwen2.5-Coder-7B model on CruxEval-O problems. For each problem, we generate CoT from both models and evaluate two key dimensions: consistency between reasoning and final answer, and information content of the reasoning content.